OpenStudy (anonymous):

how do you factor 5x^2-5x-30

OpenStudy (anonymous):

first you must simplify the whole equation by a common factor, can u see what it is?

OpenStudy (anonymous):

then it will be much easier to factorize.

OpenStudy (mertsj):

First factor out the 5

OpenStudy (anonymous):

In the equation divide 5x^2, 5x and 30 by five to factor out five.

OpenStudy (anonymous):

Then you would get the equation x^2-x-6. And you would need to factor it by finding two numbers that add up to -1, and has a sum of -6.

OpenStudy (anonymous):

find the roots of the equation

OpenStudy (mertsj):

@mivanov It's not an equation.

OpenStudy (anonymous):

i meant for her to find the roots of the equation 5x^2-5x-30=0

OpenStudy (mertsj):

But the problem said to factor the expresson. That's why she can't divide by 5. She has to factor out the 5

OpenStudy (anonymous):

Yes she has to divide by 5, in order to factor.

OpenStudy (anonymous):

Oh I think I see what you mean... so the equation would look like this 5(x^2-x-6)

OpenStudy (mertsj):

Yes. And then factor the trinomial inside the parentheses

OpenStudy (anonymous):

Yeah your right I forgot about putting the 5 outside the para(), but everything else you so the same way.

OpenStudy (whpalmer4):

@maria74123 you mean add up to -1 and have a product of -6 \[5x^2-5x-30 = (5x + a)(x+b)\] for some values of a and b. We know that 5 is prime, so one of the products will have a 5x and the other an x. \[(5x+a)(x+b) = 5x^2 + 5bx + ax + ab = 5x^2 + (5b+a)x + ab\]but we want that to have the same form as our original expression, so we can see that the following must be true:\\[5b+a = -5, ab = -30\] If we look at the factors of 30, we have 2, 3, 5, 6, 10, 15 to work with. 5(2) - 15 = -5, and 2*(-15) = -30, so our factored expression is \[(5x-15)(x+2) or 5(x-3)(x+2)\]

OpenStudy (whpalmer4):

Not necessary to factor out the 5 to factor the polynomial, but it might make factoring easier.

OpenStudy (mertsj):

It is necessary if you want to factor completely.

OpenStudy (mertsj):

Any typically, instructors want things factored completely.

OpenStudy (whpalmer4):

No, I wasn't clear - it isn't necessary to factor the 5 out before factoring the trinomial.

OpenStudy (whpalmer4):

I agree that it needs to come out somewhere for full factoring, but it doesn't have to be first.

OpenStudy (anonymous):

Yeah I always learned to factor out that way, because it is easier. But everyone does it differently.... and yes I said that wrong, I meant find the product of -6, not sum Oops lol

OpenStudy (mertsj):

@whpalmer4 Maybe not necessary but it certainly makes life easier to take out the 5 first.

OpenStudy (anonymous):

All this work, and this girl juicepop2i sn't even online anymore Haha people drive me nuts!

OpenStudy (whpalmer4):

It does, but practice factoring things where the leading coefficient != 1 makes factoring them when you can't simplify it more familiar. Practice doing more difficult things makes them easier when they can't be avoided.